Shift report 20000828 16:00-24:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Mon Aug 28 2000 - 00:02:30 EDT

  • Next message: http account: "Shift report 20000828 08:00-16:00"

    ------------------------------------------------------------
    Date: Mon Aug 28 00:02:29 2000, Shift: 16:00-24:00
    Supervisor: FV, with: FV,AJ, Bergen
    Log book#5 pages 28 - 34
    
    Shift Summary:
    ------------------------------------------------------------
    Machine was off during day (8 to about 19) for the RHIC open house
    During this period some work was done in the IR
    
    (RAS,KO,RD,ZM)
    
    i) Re-install and checkout of L1458B with the repaired
    powersupply module (RAS, ZM)
    ii) Interchange of TPC trigger cable between T1 and T2.
     The sync error rate is now acceptable (85 in two hours
      with beam this evening).
    iii) Checkout of T3,T5 with Bertran Supplies. After closeup
      it was realized there is some problem with a T5 sub module
      drawing excessive current.
    iv) Checkout of fb1 readout i.e. for WM2. 
    v)  Installation of Hall probe in D2.
    vi) connection of tpc drift velocity monitor for TPM2. HV and
    signal cables connected to HV supplies and FEHut. Await inclusion in
    data readout (Camac?)
    vii) No apparent obstacles on MRS rails (around 63 deg.)
    
    viii) Two tiles were hung from D5 behind TPM1 and cable runs
     checked out.
    ix) Called CAS watch to ensure all magnets D1-D5 can be turned
     on. RD raised concern about polarities for D2. When D5 and D2 are both at
    B-setting Ramiro expected to see same sign field for Hal probes
    but saw opposite sign . Needs investigation.
    
    Area was closed up at 17.00. 
    Beam accelerated and cogged at about 20.30 with intensities of
    B (16e09) Y (17e09) but
    
    1. ZDC coinc rates were ~1200 a factor 4 bigger than expected
       andmost single rates (BB, H1 and Yellow ZDC) a factor of 20
       higher than previous runs. The TPC anode wires were all drawing 
       high leakage currents.
    2. After must talking, as well as investigations by MCR, it
       was realized that work has gone on during the day on the
       yellow ring (warm sector 3) possible to improve vacuum but in
       effect the vacuum was up to e-6 and thus cause for the very bad
       background. It was estimated that it can take at least a day
       before this becomes good again. During this MCR had invesitgated bumps
       and other poosible causes.
       This background do not show up at other experiments, but
       causes a much faster decay of yellow beam. 
    3. I informed MCR that this beam is unusable for BRAHMS and 
       will shut experiment down
       until vacuum has improved. It should be followed up at
       tomorrws 8.30 meeting (Dana) at what time we hopefully 
       have better estimate of when vacuum is improved.
       MCR will make triplet studies 
       around 2 o'clock. Quench link interlock at 23.05.
    
    4. Magnet turned stby/off. Beam Beam HV mainframe turned 
       off.
    
    5. One data run 2457 was taken under the bad beam 
       conditions. (240K events at about 70 sec). All triggers
       were downscaled. 
    
    Thus in short an unhappy ending after the productive opening during
    the day.
    
    
    
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Mon Aug 28 2000 - 00:06:03 EDT